OpenStudy (anonymous):

j-(-8)=5 so j=13 right

Directrix (directrix):

j - (-8 )= 5 j + 8 = 5 j + 8 - 8 = 5 - 8 j = -3 ---> A different answer --> Please post your work so that we can diagnose the error or find my error.

OpenStudy (kymber):

I think we've already discussed this. If there are two negative signs, they cancel out and become positive. So it's actually J+8=5, and you would subtract the eight, making it -3.